OpenStudy (anonymous):

what is the product of (x+3)(x-3)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

x^2-9

OpenStudy (anonymous):

\[x ^{2}-9\]

OpenStudy (anonymous):

yes @krischristian is right why don't do the multiplication and find out if we are right?

OpenStudy (uri):

(x+3)(x-3) x^2-3x+3x-9 x^2-9 Ans -3x +3x gets cancelled!

OpenStudy (anonymous):

well done uri

OpenStudy (anonymous):

x^2 - 9

OpenStudy (anonymous):

(x+3)(x-3) x times x = x^2 +3-3=0 +3 times -3=-9 Answer=x^2-9 (dont include 0)

OpenStudy (anonymous):

use square - difference - product

OpenStudy (anonymous):

Since the equation you gave was unlike signs binomial, If it's like signs, it is square + sum + product